OpenStudy (anonymous):

Find sin θ and cot θ of 480°.

OpenStudy (anonymous):

so would it be 480-360 =120 then it would be 180-120 = 60 which means it's a 60 degree angle? then it would be: sin = sqrt(3)/2 and cot = sqrt(3)/3 is this correct?

OpenStudy (mertsj):

Well yes,except you forgot that it is in the second quadrant and the cot is negative in Quadrant II

OpenStudy (anonymous):

oh yea snap, lol. Thanks for the catch.

OpenStudy (mertsj):

yw